### v2

# Capacitance-based Fermion parity read-out and predicted Rabi oscillations in a Majorana nanowire

Abstract
Recent experiments have measured flux dependent capacitance at radio frequencies leading to the potential for a fast parity read-out of a Majorana qubit. In this work we argue that the quantum dot used in the capacitance measurement can be reasonably approximated by a non-interacting weakly coupled orbital. We then predict the measured flux and parity dependent capacitance for several parameter regimes of the disordered Majorana nanowire model that are both topological and trivial. Following this we study how such a fast capacitance read-out can be used to characterize the quantum coherence of a Majorana nanowire-based qubit using Rabi oscillations. We additionally show that such measurements, if made possible by coherent inter-wire tunneling, would provide a valuable way of characterizing the low energy states in the frequency domain.
The topologically protected degree of freedom associated with a Majorana system is the Fermion parity of the pair of Majoranas at the end of a wire. This Fermion parity is equivalent to the occupancy of a low-energy Andreev bound state (ABS), seen in Fig. 1, that is formed from the pair of Majorana zero modes (MZMs). The addition of tunneling between the ends results in a closed loop [1] with the ABS moved to finite energy, which is shown on the left half of Fig. 1. In the case of an ideal topological wire, the quantum information (i.e. Fermion parity) of the pair of Majorana modes transforms into the occupancy of this finite energy bound state. Any scheme to measure the occupancy of this finite energy bound state by measuring inductance [2], conductance [3, 4] or by a capacitance measurement [5, 6] constitutes a read-out of the Majorana qubit. Such a capacitance measurement-based parity readout in a Majorana nanowire [2, 7] has recently been accomplished using a fast RF circuit [8]. A capacitance measurement contains contributions other than that from the occupied bound state. Therefore only the change in the capacitance resulting from a $\Phi_{0}$ change in the magnetic flux $\Phi$ through the loop shown in Fig. 1 is used to determine the bound state occupancy. This interferometric flux dependence of the capacitance is related to the fractional Josephson effect [9, 2] as well as non-local teleportation [10].
Figure 1: Schematic for Majorana nanowire device, similar to recent experiment [8], containing a pair of Majorana nanowires to measure Rabi oscillation. Both Majorana nanowires are assumed to be tuned into a parameter regime to support a low energy ABS, ideally formed from the pair of MBSs in each wire. The qubit is formed from the occupancy of the occupancy of $ABS_{L,R}$, which are expected to be at zero energy in the absence of tunneling (i.e. $\lambda=0$). The Majorana wire on the left containing $ABS_{L}$ is attached to a QD at both ends, forming a loop that encloses magnetic flux $\Phi$ that is controlled by a small perpendicular field. Measuring the flux dependent capacitance of the QD allows a fast read-out of the occupancy of $ABS_{L}$. Turning on tunneling $\lambda\neq 0$ splits the $ABS_{L,R}$ leading to Rabi oscillations in the occupancy of $ABS_{L}$ as a function of time that can establish the coherence of MBS qubits.
However, such an interferometric measurement of the bound state occupancy is affected by non-idealities such as the short system size relative to the effective coherence length [4]. This can be a problem even in nominally long wires such as the $3\,\mu m$ long wires used in recent experiments [8] because of sub-gap states generated by disorder [11, 12]. This makes it critical to investigate such a capacitance measurement using a general model of a disordered Majorana nanowire, where one is not limited to the low-energy subspace of the lowest Andreev states in the wire [8]. Given the demonstration of fast parity read-out [8] in a class of cleaner devices that have shown improved potential to realize MZMs, one may hope to demonstrate a higher level of coherence than was obtained in past Majorana nanowire devices [13], thus paving the way toward a Majorana qubit.
In this work, we compute the ac capacitance of a Josephson junction formed by connecting the ends of a disordered Majorana nanowire shown in Fig. 1. The capacitance is equivalent to the response of the charge in the quantum dot (QD) in Fig. 1 in response to modulating the potential of the QD. This is calculated using linear response of the Josephson junction formed through the QD, using a formalism that is similar to the calculation of the Josephson inductance [14]. We study this response for the disordered Majorana nanowire model in a variety of regimes including the Andreev bound state (ABS) and topological patch regime [11, 12]. The fast parity read out can be used to measure Rabi oscillations. We then compute the coherence, frequency and amplitude of the Rabi oscillation in topological qubits created by tunnel coupling pairs of Majorana wires in the corresponding parameter regimes.
The left Majorana nanowire in Fig. 1 contains a QD tunnel-coupled to both ends, enclosing a flux $\Phi$. The QD enables tunneling via elastic co-tunneling [15], including cross-Andreev processes [16, 17, 10], which effectively act as a direct tunneling amplitude between the nanowire ends [8]. Multi-particle excitations [15] generated in the co-tunneling process are virtual and only renormalize the tunneling amplitude. This tunneling is modulated by the gate voltage $V_{QD}$, also used for capacitance measurements. We model the QD as a single non-interacting orbital with energy level $V_{QD}$, adjusting tunneling to match capacitance data. This allows parity measurement using a non-interacting BdG Hamiltonian for the Majorana nanowire [18] (see Sec. $S_{1}$ in SI for details). The magnetic flux $\phi$ is in units where the Hamiltonian periodicity matches the flux quantum $hc/e=2$. Topological stability ($T_{S}$) is defined as the class D topological invariant [1] multiplied by the ratio of the lowest eigenvalues of the Hamiltonian (without the QD) at $\phi=0$ and $\phi=1$, helping assess wire topology and length effects. The QD hopping efficiency is set to $\lambda_{QD}=0.03$ to match energy splittings in the tens of mK, consistent with experiments. For disordered wires, the disorder potential follows a Gaussian distribution with correlation length $\xi_{dis}\sim 30nm$ and root mean square amplitude $V_{RMS}$.
The quantum capacitance measured from this gate (up to a lever arm) is given by $C=-d^{2}E_{tot}/dV_{QD}^{2}$, where $E_{tot}=\sum_{E_{m}<0}E_{m}$ is the total energy of the system. The finite frequency generalization of this quantum capacitance can be obtained for the above nanowire Hamiltonian using linear response theory as:
$C(\omega)=\Gamma_{C}\sum_{E_{n}<0,E_{m}>0}\frac{|\langle E_{n}|\tau_{z}\delta_{r,0}|E_{m}\rangle|^{2}(E_{n}-E_{m})}{(E_{n}-E_{m})^{2}-\omega^{2}},$ (1)
where $\Gamma_{C}=2\alpha^{2}e^{2}/k_{B}=1.9\,fF-K$ is a constant that converts the capacitance $C(\omega)$ to femtofarads (fF) when the energies $E_{n}$ are in units of Kelvin (K) after adding a lever arm $\alpha=2^{-1/2}$ that would be relevant to recent experiment [8]. In the above |$\ket{E_{n}}$ are eigenstates of the BdG Hamiltonian of the system, and $\tau_{z}$ is the matrix coupling to charge density, both written in the Nambu basis (see SI Sec.$S_{1}$ for details). The read-out scheme used in the recent experiments [8] involves studying the dependence of $C$ on the magnetic flux $\phi$ as well as the parity of the wire, which can be changed by switching the occupancy of the state of the smallest energy states from 00 to $1$ and vice-versa.
Figure 2: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire without disorder at $\mu=0\,K$, $B_{x}=0.28\,T$ and $V_{QD}=0.5\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability [1] (see Sec. $S_{1}$ of SI for details) $T_{S}=-0.99$ and MZM splitting for the isolated wire $E_{M}=30\,\mu K$. As expected [8] we see a shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.49$. The parameters for the phase are $B_{x}=0.4\,T$, $\mu=5.5\,K$, $V_{QD}=5.5\,K$ and $L=3\,\mu m$. The low energy states are ABSs generated by a potential (see text for details) with parameters $V_{abs}=4.5\,K$ and $L_{abs}=200\,nm$.
Consistent with the results of the experiment [8], we find that the capacitance read-out for an ideal Majorana nanowire shows an $hc/2e$ shift in flux between even and odd parity states. Fig. 2(a) shows the lowest energy (positive) state, which is composed of a pair of MZM wave-functions that are localized at the end for the optimal parameters (see caption) of a clean nanowire with the next excited state being delocalized. The individual MZM wave-functions can be constructed by considering superpositions of the lowest positive energy state, referred to as ground state in Fig. 2, with its particle-hole conjugate state with negative energy. Introducing tunneling between the ends of the nanowire through a QD to form a loop with enclosed flux splits the end MZMs leading to a pair of flux dependent energy levels shown in Fig. 2(b). The two levels represent the energies of different Fermion parity states of the nanowire. It is clear that these levels interchange energies when the flux changes by $\delta\Phi=hc/2e$ as happens in the fractional Josephson effect [1, 9, 3]. Fig. 2(c) shows the corresponding capacitance, which reflects the response of each energy state to a variation in the tunneling strength which is modulated by the potential of the QD. The capacitance shows a similar characteristic to that of the energy. The flux and parity dependence of the predicted capacitance allows, in the ideal case, a read-out of the Fermion parity of the Majorana nanowire through essentially determining the level occupancy of the split MZMs.
While capacitance-based parity readout works well for ideal Majorana nanowires, its effectiveness in regimes [19] with trivial zero-bias peaks from disorder or end potentials remains uncertain. In Fig. 2(d–f), we examine cases where zero modes originate from non-topological near-zero-energy ABSs, often induced by end quantum dot potentials [20, 21, 22, 23, 24]. Fig. 2(d) shows such ABSs forming due to an end quantum well potential (see SI Sec. $S_{1}$). Despite their low energy, Fig. 2(e) indicates that ABS spectra exhibit flux modulation smaller than the average energy, confirming the wire’s non-topological nature. Capacitance readout in Fig. 2(e) reveals a clear parity difference, but the flux dependence is dominated by an $hc/2e$ periodic modulation that flips sign between parities, unlike the ideal Majorana case where the $hc/e$ component maintains its sign. This aligns with recent capacitance analyses based on a minimal low-energy ABS model [8].
Figure 3: (a) Wave-function of the lowest energy eigenstate, (b) energy in K and (c) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=3\,\mu m$ Majorana nanowire in the topological patch regime [11] at $\mu=0.2\,K$ and $B_{x}=0.42\,T$, $V_{QD}=0.2\,K$, disorder strength $V_{RMS}=7.5\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability [1] (see Sec. $S_{1}$ of SI for details) $T_{S}=-0.621$. As expected [8] we see an approximate shift of $\Delta\phi=1$ between the even and odd parity sectors. The panels (d,e,f) are the corresponding result for the Majorana nanowire in a non-topological phase with $T_{S}=0.13$. The parameters for the phase are $B_{x}=0.56\,T$, $\mu=2.0\,K$, $V_{QD}=2.0\,K$ $V_{RMS}=15\,K$ and $L=3\,\mu m$.
Three-terminal measurements of tunnel barrier dependence can rule out local zero-energy ABSs [25]. However, low-energy states from intermediate disorder, known as topological patches [11, 12], may explain experimentally observed topological regions [25]. Fig. 3(a) shows the lowest-energy eigenstate in a weakly disordered Majorana nanowire, where despite a negative topological stability ($T_{S}$), the ABS wavefunction lacks significant end localization, distinguishing it from true MZMs. Still, Fig. 3(b) reveals an $hc/2e$ flux shift between parity sectors, resembling the MZM spectrum in Fig. 2(b). Similarly, capacitance in this regime (Fig. 3(c)) exhibits a notable $hc/e$ component, flipping sign like in the topological case, though with asymmetric oscillation amplitudes, consistent with some experimental samples [8]. Topological stability in the patch regime [11] is highly disorder-dependent. Stronger disorder (Fig. 3(d)) leads to shorter localization compared to Fig. 3(a). Despite a trivial $T_{S}$ ($T_{S}>0$), Fig. 3(e–f) still shows substantial $hc/e$ components, resembling the topological case in the upper panels.
Figure 4: Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with a scaling factor of $0.76$ for the top panels and $0.01$ panel (c) and $0.3$ for panel (d)) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. The left panels (right panels) refer to the case for even (odd) total Fermion parity respectively of the system shown in Fig. 1. The top panels (a,b) correspond to parameters of Fig. 2(a,b,c) with inter-wire parity decoherence rate $(T_{1}^{\prime})^{-1}\sim 38\,Hz$ while the bottom one corresponds to parameters Fig. 2(d,e,f) with $(T_{1}^{\prime})^{-1}\sim 0.01\,GHz$.
To characterize the coherence of a nanowire-based qubit, we propose using Rabi oscillations, in the two Majorana wire set-up coupled by tunneling shown in Fig. 1. The Hamiltonian of the system is
$H_{tot}=H_{L}+H_{R}$
$+\lambda(\tau)[(t+it_{SO}\sigma_{y})\tau_{z}\ket{r=N,L}\bra{r=1,R}+h.c],$ (2)
where $H_{L,R}$ are given by Eq. 3 except that the sites are restricted to $r\neq 0$ together with $\lambda_{QD}=0$ to remove the QD and $\lambda(\tau)$ is the dimensionless tunneling strength between the wires. Note that this set-up differs fundamentally from previous proposals for studying coherence in nanowire qubits [26], which requires inter-wire parity measurements. which requires inter-wire parity measurements. Instead, the set-up in Fig. 1, similar to previous experiments [13], is based on intra-wire parity measurements together with inter-wire tunneling that is set to a non-zero value $\lambda(\tau>0)=\lambda$ after initializing with $\lambda(\tau=0)=0$. Restricting the dynamics to the lowest two positive energy state (one from each nanowire) and considering a specific total Fermion parity sector allows one to map the dynamics using the Bloch-Redfield master equation of a standard two-level spin system [27] that is characterized by a Rabi oscillation frequency $\Omega$, relaxation time $T_{1}$, dephasing time $T_{2}$ as well as an amplitude for the Rabi oscillation of the Fermion parity of the left wire (see SI Sec. $S_{2}$ for details). The decoherence is generated by bath couplings that couple to the gate voltage in each wire with magnitude $\gamma_{L,R}\sim 80\,mK$ and are chosen to reproduce conventional charge qubit dephasing [28] (see SI Sec. $S_{3}$ for details). Inter-wire parity or capacitance measurements, which are necessary to perform topological operations [6], allow one to estimate coherence of the qubits [26] without inter-wire tunneling. The estimate of dephasing rates in this case takes the form of a decoherence time $T_{1}^{\prime}$ for the inter-wire fermion parity. Different choices of ends for the inter-wire fermion parity measurement $X$ and $Y$ are expected to generate different superpositions that differ by a phase shift $\delta\phi$ that can also characterize a topological qubit (see SI Sec. $S_{4}$ for details).
We first consider the coherence properties of a qubit constructed out of a pair of ideal Majorana nanowires whose capacitive read-out properties have been described in Fig. 2(a-c). We find, in Fig. 4(a,b), that the Rabi oscillation frequency $\Omega$ varies linearly with the tunnel coupling parameter $\lambda$ between the wires. The panels $a$ and $b$ differ by the total Fermion parity of the wire system. In contrast, the amplitude of the Rabi oscillation, which is essentially perfect (i.e. equals $0.5$) is independent of $\lambda$. The relaxation time $T_{1}$ and the dephasing time $T_{2}$ are, as expected, also essentially infinite as seen in these panels. The inter-wire parity decoherence rate $(T_{1}^{\prime})^{-1}$ is also quite small on this scale. In contrast, the behavior of the Rabi oscillations for the ABS case corresponding to Fig. 2(d-f), which is shown in Fig. 4(c,d), is quite different and shows a Rabi oscillation frequency that varies relatively weakly with tunnel coupling $\lambda$. Instead the amplitude of the Rabi oscillation is what vanishes as $\lambda$ decreases. Additionally, the Rabi oscillation amplitude is strongly suppressed in the even fermion parity sector. This suggests that the Rabi oscillation originates from a complicated interplay between inter-wire states and intra-wire ABSs. Surprisingly (though consistent with suggestions from prior work [29]), we find that the relaxation and dephasing times $T_{1,2}$ and $T_{1}^{\prime}$ are still extremely long for the ABS based qubits, though the phase shift $\delta\phi$ is different from the topological value. This clearly implies that the superconductivity by itself not just topology can play a strong role in screening the charge noise that generates dephasing in non-superconducting charge qubits, which should be readily observable experimentally.
Figure 5: Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.3$ for the upper panels, $0.1$ for (c) and $1.1$ for (d)) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. The left panels (right panels) refer to the case for even (odd) total Fermion parity respectively of the system shown in Fig. 1. The top panels (a,b) correspond to parameters of Fig. 3(a,b,c) with inter-wire parity Rabi oscillation dephasing rate $(T_{2}^{\prime})^{-1}\sim 0.02\,GHz$ while the bottom one corresponds to parameters Fig. 3(d,e,f) with $(T_{2}^{\prime})^{-1}\sim 0.42\,GHz$.
Fig. 5 shows Rabi oscillations in weakly disordered Majorana nanowires in the topological patch regime corresponding to the parameters in Fig. 3. The results in Figs. 5(a,b) show that the topological regime (i.e. $T_{S}<0$) corresponding to Figs. 3(a,b) show relatively coherent Rabi oscillations with relaxation and dephasing rates $T_{1,2}^{-1}$ that are a fraction of the Rabi oscillation frequency provided the dimensionless tunnel coupling $\lambda$ can be increased to a reasonable fraction. This, of course, is much less coherent compared to the disorder free MZM model shown in Fig. 4(a,b). The crucial requirement of being in the weak disorder limit becomes apparent from the lack of Rabi oscillations with larger disorder seen in Fig. 5(c,d), which shows that the Rabi frequency is a small fraction of both the relaxation and dephasing rate $T_{1,2}^{-1}$.
Comparing the results for our simulated flux and Fermion parity dependence of the capacitance for the ideal MZM case (i.e. Fig. 2(a,b,c)) to the other cases suggest that the capacitance in the topological case is characterized by an $hc/e-$periodic flux dependence together with an exact $hc/2e$ phase shift between the different parity sectors. Indeed, as shown in the SI Sec. $S_{5}$, both by explicit calculation of low energy ABSs as well as the fractional Josephson effect [9, 1, 2, 3] that this is a definitive signature of MZMs. A gapped trivial phase shows an $hc/2e-$periodic flux dependence similar to Fig. 2(d,e,f). The topological patch regime Fig. 3, while showing a near $hc/2e$ phase shift between the different parity sectors, also shows a change in average as well as amplitude that are not allowed in the topological phase. The recent capacitance measurements in the Majorana nanowire [8], do not distinguish between MZM and the topological patch regime because of an ambiguous flux periodicity.
Furthermore, our analysis establishes that the samples studied in Ref. [8] should manifest Rabi oscillations.In fact, we find in Figs. 35 that the coherence of the Rabi oscillation matches the inter-wire parity coherence $T_{1}^{\prime}$, which in turn is proportional to the gate charge noise. This suggests that in the case of devices containing wires from different categories in Figs. 35 the coherence would be limited by the lower quality nanowire. The combination of flux dependence capacitance, coherent tunneling Rabi oscillation as well as inter-wire parity Rabi oscillations can indicate thus providing a direct qubit diagnostic for Majorana nanowires.
Acknowledgements.
JS acknowledges valuable discussions with Mike Manfra, Bela Bauer and using KWANT code from Bas Nijoholt’s github (https://github.com/basnijholt/majorana-nanowire-conductance). This work is supported by the Laboratory for Physical Sciences and by the Joint Quantum Institute co-directors fund.
Appendix A Supplementary Information
A.1 $S_{1}$: Details of Majorana nanowire model with quantum well and disorder
THe Hamiltonian for the system is written as:
$H=\sum_{r\leq N}\{(-2t-\mu+V_{dis}(r))\tau_{z}+\tau_{x}\Delta_{0}+V_{Z}\sigma_{z}\}\ket{r}\bra{r}$ |
$|\displaystyle+[\sum_{0<r<N}(t+it_{SO}\sigma_{y})\tau_{z}\ket{r+1}\bra{r}$ |
$+\lambda_{QD}(t+it_{SO}\sigma_{y})(\ket{r=N}-e^{i\pi\phi\tau_{z}}\ket{r=1})\bra{r=0}+h.c],$ (3)
where site $r=0$ represents the QD and $\lambda_{QD}$ is the relative weakening of the tunnel barrier between QD and the nanowire. The Hilbert space for the above Hamiltonian, contains a $4$ dimensional Nambu space at each site $r$, with $\sigma_{x,y,z}(\tau_{x,y,z})$ being Pauli matrices in spin (particle-hole) space. The Rashba spin-orbit coupling with strength $\alpha_{R}$ is perpendicular to the wire and the magnetic field $B_{x}$ is applied along the nanowire so as to generate with Zeeman term of strength $V_{z}=(g/2)\mu_{B}B_{x}$ (with the factor chosen to be $g\mu_{B}=15\,K/T$ based on recent experiments [25]). Based on recent experimental characterization of high quality Majorana nanowire devices [25], we choose $t_{0}=253.8\,K$ (corresponding to an effective mass $m^{*}=0.034$ for InAs and lattice parameter $a=10.1\,nm$ ) $t_{SO}=8.59K$ ( corresponding to Rashba coupling $\alpha=8\,meV-nm$), and the superconducting coupling is $\Delta_{0}=1.5\,K$ corresponding to the low frequency limit of the superconducting self-energy [30]. The hopping efficiency $\lambda_{QD}$ through the QD is chosen to be $\lambda_{QD}=0.03$ so as to obtain energy splittings in the range of severeal tens of mK. In the above $V_{dis}(r>0)$ is the disorder potential for the Majorana nanowire which is taken to be Gaussian distributed with a correlation length of $\xi_{dis}\sim 30nm$ with an RMS amplitude $V_{RMS}$, while $V_{dis}(r=0)=V_{QD}$ is the QD potential set by a gate. In cases where an end potential generated ABS is present, the zero-energy end ABSs are generated by adding an end square quantum-well potential of width $L_{abs}=200\,nm$ and depth $V_{abs}=4.5\,K$ into $V_{dis}(r)$ in Eq. 3. The profile of the quantum well is somewhat arbitrarily chosen based on previous work [20, 21, 22, 23, 24] so as to produce low energy localized ABSs as seen in Fig. 2(c).
The topological stability $T_{S}$ used to characterize the topological nature of the wire is closely related to the topological invariant of the Majorana nanowire defined by Kitaev [1]. The topological invariant or Majorana number was defined by Kitaev for the Hamiltonian in Eq. 3, which has periodic/antiperiodic boundary conditions as the ground state fermion parity of $H$. The ground state fermion parity of $H$ can be written as $P(\lambda_{QD})=\textrm{sign}(Pffafian(\sigma_{y}\tau_{y}H(\lambda_{QD})))$. The topological invariant was defined by Kitaev [1] as the ratio of the Fermion parity between periodic and anti-periodic boundary conditions $P_{top}=P(\lambda_{QD}=1)/P(\lambda_{QD}=-1)$. This topological invariant provides no insight into the proximity to the topological transition. The Pffafian in $P$ changes sign when the smallest eigenvalue of $H(\lambda_{QD}=\pm 1)$ crosses 00. For this reason we define the topological stability
$T_{S}=P_{top}\textrm{min}(E_{g}(\lambda_{QD}=\pm 1))/\textrm{max}(E_{g}(\lambda_{QD}=\pm 1)),$ (4)
where $E_{g}(\lambda_{QD})$ is the gap of the Hamiltonian $H(\lambda_{QD})$.
A.2 $S_{2}$: Technical details of Rabi oscillation calculation
The Rabi oscillation is observed by measuring the Fermion parity, which in the low energy sector is given by $F_{P,L}=\ket{\psi_{1,L}}\bra{\psi_{1,L}}-\ket{\tilde{\psi}_{1,L}}\bra{\tilde{\psi}_{1,L}}$ | of $H_{L}$. Here |$\ket{\psi_{1,L}}$ and |$\ket{\tilde{\psi}_{1,L}}$ are the smallest positive and negative energy eigenstates of the BdG Hamiltonian $H_{L}$ at $\lambda=0$. Particle-hole symmetry relates these two states so that the occupancy of one state guarantees that the other state is unoccupied. In this work, we make the simplifying approximation that following the turn-on of the tunneling for $\tau>0$, only the lowest two positive energy states $\phi_{1,2}$ (and their particle-hole conjugates $\tilde{\phi}_{1,2}$) with energies $E_{1,2}$ can be occupied in a way that is constrained by the total Fermion parity of the system being even or odd. Particle-hole symmetry means that the occupancy of the other states $\tilde{\phi}_{1,2}$ is determined by the occupancy of their particle-hole conjugate. Therefore, the Rabi oscillation for $\tau>0$ can be considered to either be an oscillation between the single particle levels $\phi_{1}$ and $\phi_{2}$ for the case of odd total parity and $\phi_{2}$ and $\tilde{\phi}_{1}$ for the case where the total parity is even. This allows one to describe each parity sector by a two-level system subject to bath coupling operators $S_{L}=\tau_{z}\sum_{r}\ket{r,L}\bra{r,L}$ | and $S_{R}=\tau_{z}\sum_{r}\ket{r,R}\bra{r,R}$ | that represent uniform potential coupling to the left and right wires respectively. Assuming that the noise on the two wires is uncorrelated, the Linblad form of the Bloch-Redfield master equation [27] is written as:
$\dot{\rho}=-i[H_{tot},\rho]$
$+\sum_{a=L,R;\omega}\gamma_{a}(\omega)\left[S_{\alpha}(\omega)\rho(\tau)S_{\alpha}^{\dagger}(\omega)-\frac{1}{2}\{S_{\alpha}(\omega)S_{\alpha}^{\dagger}(\omega),\rho(\tau)\}\right],$ (5)
where $S_{\alpha}(\omega)=\sum_{\omega=\omega_{a}-\omega_{b}}\Pi_{a}S_{\alpha}\Pi_{b}$ and $\Pi_{a}$ are projection into eigenstate |$\ket{\omega_{a}}$ of $H_{tot}$ with energy $\omega_{a}$. Note that $\rho(t)$ is the density matrix containing only the two single-particle levels of $H_{tot}$ appropriate for the relevant parity sector. We choose the limit of large bath temperature so that $\gamma_{\alpha}(\omega)/\gamma_{\alpha}(-\omega)=e^{-\omega/T_{bath}}\approx 1$. The parameters $\gamma_{L,R}$ are noise powers that are set to $\gamma_{L}=80$ mK and $\gamma_{R}=84$ mK consistent with decoherence rates for non-superconducting charge qubits. The charge noise is assumed to act uniformly on the wire according to $Q_{\alpha}=\tau_{z}\sum_{r}\ket{r,\alpha}\bra{r,\alpha}$ |, which in the basis of the Rabi oscillator eigenstates is $S_{\alpha,a,b}=\langle\psi_{a}|Q_{\alpha}|\psi_{b}\rangle$.
The Rabi oscillation in the low-energy case for a fixed Hamiltonian and two low energy state maps explicitly to the spin1/212-1/2- 1 / 2 (i.e. two level system). In the basis of eigenstates |$\ket{\omega_{\pm}}$ the Hamiltonian $H_{tot}=\epsilon\sigma_{z}$ and the projector operators are $\Pi_{\pm}=\frac{1\pm\sigma_{z}}{2}$. The solution $\rho(\tau)$ of Eq. 5 is a Hermitean matrix with unit trace and therefore can be written in Bloch form as $\rho(\tau)=\mathbf{1}/2+\mathbf{r}(\tau)\cdot\mathbf{\sigma}$, where $\mathbf{r}(\tau)$ is a real $3-$ vector. The Bloch-Redfield equations can be written as $d\mathbf{r}/d\tau=M\mathbf{r}(\tau)$, where $M$ is a real $3\times 3$ matrix. One of the eigenvalues of $M$ is purely real $T_{1}^{-1}$, while the other two are a complex conjugate pair $T_{2}^{-1}+i\Omega$. The corresponding eigenvector $r_{0}$ of $M$ defines the Rabi oscillation of the observable $O$ given by
$\textrm{amplitude}=|Tr[\mathbf{r}_{0}\cdot\mathbf{\sigma}\tilde{O}]|,$ (6)
where $\tilde{O}$ defined by the matrix elements $\tilde{O}_{ab}=N\langle\omega_{a}|O|\omega_{b}\rangle$ are the matrix elements of the observable $O$ in the low energy subspace and $N$ is a normalization factor chosen so that $Tr[\tilde{O}^{2}]=1$. The operator $O$, in these nanowire devices, which appear in the Hamiltonian as a modulated tunneling term $\lambda(\tau)O$ can dominate the Hamiltonian in the low-energy sector. In this case, the measurement of the responses $\partial^{2}H/\partial\lambda^{2}\propto O$ constitutes a measurement of the operator $O$. For the Rabi oscillation presented in the paper, $O$ is the end-end tunneling in a single wire i.e.
$O_{Z}=\tau_{z}[e^{i\phi\tau_{z}}\ket{r=0,L}\bra{r=N,L}+h.c]$ (7)
applied to the decoupled wires, where $\phi_{Z}$ is the magnetic flux-induced phase in the tunneling. In the limit that the wire has a single (pair if one looks at the BdG spectrum) low energy state in the wire, the fermion parity $F_{P}$ (or equivalently the occupation of this state) is the only degree of freedom so that the operator $O_{Z}=\bar{O}F_{P}$ when projected on to the low energy space becomes proportional to the Fermion parity (apart from a shift) discussed in the main text.
Figure 6: (a) Energy in K and (b) capacitance for the occupied and unoccupied state as a function of magnetic flux $\phi$ (in units where $hc/e=2$) for an $L=500\,nm$ Majorana nanowire in the topological patch regime [11] at $\mu=-1.62\,K$ and $B_{x}=0.28\,T$, $V_{QD}=0.0\,K$, $V_{RMS}=0\,K$. The wave-functions for this ideal case show clear presence of end MZMs with a topological stability (defined below Eq. 3) $T_{S}=-0.06$. (c,d) Rabi oscillation parameters relaxation rate $T_{1}^{-1}$, dephasing rate $T_{2}^{-1}$, Rabi-oscillation frequency $\Omega$ and amplitude (dimensionless with scale of $0.2$ for the upper panels and $14$ for the lower panels) of Fermion parity oscillation as a function of inter-wire coupling $\lambda$. Panels (c) and (d) refer to Rabi oscillations in the even and odd total Fermion parity respectively of the system.
A.3 $S_{3}$: Rabi oscillation calibration
We calibrate the noise parameters for the master equation by calculating the Rabi oscillations in a nanowire configuration similar to the set-up discussed in the main text, other than that superconductivity is removed. Without superconductivity the Fermion parity qubit simply because a conventional charge qubit [28]. The result of this non-superconducting case that is used to set the value of $\gamma_{L,R}$ is shown in Fig. 6.
A.4 $S_{4}$: Inter-wire parity Rabi oscillations
An alternative way to test coherence is to consider the case where the tunneling introduced is inter-wire so that
$O_{X}=\tau_{z}[e^{i\phi\tau_{z}}\ket{r=N,L}\bra{r=0,R}+h.c]$ (8)
$O_{Y}=\tau_{z}[e^{i\phi\tau_{z}}\ket{r=0,L}\bra{r=0,R}+h.c].$ (9)
Again, in the case where there is a single low energy state in each wire, $O_{X,Y}$ flip the occupancy of the state in each wire and therefore anti-commute with the Fermion parity of the left wire $F_{P}$. This means that $O_{X,Y}$ also anticommute with $O_{Z}$. This is part of the consistency requirement for an ideal Majorana representation where $O_{Z}=i\cos{\phi}\gamma_{LL}\gamma_{R,L}$, $O_{X}=i\cos{\phi}\gamma_{LL}\gamma_{LR}$ and $O_{Y}=i\cos{\phi}\gamma_{RL}\gamma_{RR}$ would form three anti-commuting Pauli matrices. In the general case of a single low-energy level in each wire, which encompasses the patch regime as well as the single ABs regime, the anti-commutation between $O_{X}$ and $O_{Y}$ is not obvious.
To understand the states of the qubit at fixed fermion parity we define the $O_{Z}$ (or $z$-basis) of the qubit as a fermion being on the left nanowire |$\ket{L}$ or on the right-nanowire |$\ket{R}$. The even or odd total fermion parity is then simply about choosing a reference ground state with even or odd fermion parity. The two states |$\ket{L,R}$ are the two eigenstates of $O_{Z}$ or equivalently $F_{P,L}$ (i.e. fermion parity of the left wire). Measurement of the inter-wire Fermion parity $O_{X,Y}$, which anticommutes with $F_{P,L}$ then leads to a superposition state |$\ket{L}\pm e^{\pm i\phi_{X,Y}}\ket{R}$ where $\phi_{X}$ and $\phi_{Y}$ are phases associated with the measurement $O_{X},O_{Y}$. A topological qubit would result in a phase-shift $\delta\phi=|\phi_{X}-\phi_{Y}|-\pi/2\sim 0$ between the measurements $O_{X}$ and $O_{Y}$ that would be necessary for the implementation of Pauli gates in the Clifford algebra [6]. The calculation of such a phase shift, which can be estimated from measuring the correlation between $X$ and $Y$ on the qubit, provides another comparison to a topological qubit. We find, however, that all the cases we studied except the ABS wire (i.e. Fig. 2(d,e,f)) satisfied this topological criterion. Experimentally, testing this phase shift condition would amount to testing the orthogonality of $O_{X}$ and $O_{Y}$ i.e. whether an $O_{X}$ eigenstate is a completely mixed $O_{Y}$ state.
Measurement of inter-wire tunneling allows a few more tests. Specifically, in this case, one can obtain Rabi oscillations even without the interwire tunneling. The operators $S_{\alpha}$ are diagonal in eigenstates, which are localized to $a,b=L,R$. Therefore $S_{\alpha}(\omega)$ is only non-zero at $\omega=0$ and are proportional to projection matrices i.e. $S_{\alpha}(0)=\Pi_{\alpha}D_{\alpha}$, where $D_{\alpha}=S_{\alpha,\alpha\alpha}=\langle\omega_{\alpha}|\tau_{z}|\omega_{\alpha}\rangle$. The diagonal elements of $\rho$ then become conserved and the master equation reduces to pure dephasing
$\dot{\rho}_{12}=-[i(\omega_{L}-\omega_{R})-\frac{(\gamma_{L}D_{L}^{2}+\gamma_{R}D_{R}^{2})}{2}]\rho_{12}.$ (10)
The dephasing rate in this case is simply $(T_{1}^{\prime})^{-1}=(\gamma_{L}D_{L}^{2}+\gamma_{R}D_{R}^{2})/2$, which essentially corresponds to the response of the energy level to charge noise.
A.5 $S_{5}$: Characterizing MZM from capacitance
The fractional Josephson effect [9, 1, 2, 3] in the topological superconducting phase can be used to argue for a characteristic flux dependence of the capacitance. To understand this, note that for a ”long” superconducting wire, which is longer than the coherence length, the introduction of a phase slip in the middle of the superconducting wire is equivalent to an $hc/2e$ phase shift. In the trivial phase, this cannot have an effect on the capacitance, which must then be $hc/2e-$ periodic. In contrast, in the topologically non-trivial phase, such a phase slip must be accompanied by a change in fermion parity. Thus, the capacitance in the topological case must be characterized by an $hc/2e-$ phase shift between the two fermion parity sectors, which must otherwise be $hc/e-$periodic.
To check this result explicitly, consider a low energy ABS or pair of MZMs in a wire, which can be represented by a fermion operator $c=(\gamma_{R}+i\gamma_{L})2^{-1/2}$. Let us represent the low energy fermion operator on the site $N=0$ by $d$. Introducing the flux dependent phase $\phi$ (in units so that $hc/2e=2\pi$) on the right tunnel junction:
$H=\delta c^{\dagger}c+2^{1/2}t_{R}e^{i\phi/2}(\gamma_{R}+\zeta_{R}\gamma_{L})d$
$+2^{-1/2}t_{L}d(\gamma_{L}+\zeta_{L}\gamma_{R})+h.c.,$ (11)
where $t_{L,R}>0$ are hopping amplitudes through the left and the right ends of the wire, $\zeta_{L,R}$ represent the leakage amplitudes of the MZMs across the wire and $\delta$ is the MZM splitting. Writing the Hamiltonian in terms of the fermion operator c, we get
$H=\delta c^{\dagger}c+d[t_{R}e^{i\phi/2}\{(1+i\zeta_{R})c^{\dagger}+(1-i\zeta_{R})c\}$
$+t_{L}\{(i+\zeta_{L})c^{\dagger}+(-i+\zeta_{L})c\}]+h.c.$ (12)
Note that choosing $\zeta_{R,L}$ to be complex allows $H$ to contain arbitrary complex normal and Andreev tunneling amplitude at each junction.
The two terms separate naturally by fermion parity sector. In the even fermion parity sector $c^{\dagger}c+d^{\dagger}d=0,2$, the $cd^{\dagger}=c^{\dagger}d=0$ so that writing $d^{\dagger}d=c^{\dagger}c=(1+\sigma_{z})/2$ and $c^{\dagger}d^{\dagger}=\sigma_{+}$ we get
$H_{even}(\phi)=\delta\sigma_{z}/2$
+$+\{t_{R}e^{i\phi/2}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})\}\sigma_{-}+h.c.$ (13)
In the odd fermion parity sector $c^{\dagger}c+d^{\dagger}d=1$, the $c^{\dagger}d^{\dagger}=cd=0$ so that writing $21-d^{\dagger}d=c^{\dagger}c=(1+\sigma_{z})/$ 2 and $dc^{\dagger}=\sigma_{+}$ we get
$H_{odd}(\phi)=\delta\sigma_{z}/2$
+$+\{t_{R}e^{i\phi/2}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})\}\sigma_{+}+h.c.$ (14)
Writing the Hamiltonian in each case as $H=(\delta/2)\sigma_{z}+A\sigma_{+}+A^{*}\sigma_{-}$ we can write the capacitance as
$C=\partial^{2}E/\partial\delta^{2}=\frac{2|A|^{2}}{(\delta^{2}+4|A|^{2})^{3/2}},$ (15)
where the amplitude for the even and odd case is
$|A|^{2}$ $=|t_{R}e^{i\phi/2}(1-i\zeta_{R})+t_{L}(-i+\zeta_{L})|^{2}\textrm{for even parity}$
$=|t_{R}e^{i\phi/2}(1+i\zeta_{R})+t_{L}(i+\zeta_{L})|^{2}\textrm{ for odd parity}.$ (16)
The $\phi-$dependence of $|A|^{2}$ in both cases can be written in terms of three parameters $P,Q$ and $R$ as $|A|^{2}=P+Q\cos{\phi/2}+R\sin{\phi/2}$ ). In principle, Eqn. 15 together with this parametrization allows a fit for $\delta,P,Q$ and $R$ from the flux dependence of $C(\phi)$ over one period. The parameters $P,Q,R$ for each parity sector combine to form six parameters, which can be combined using Eq. 16 to determine $t_{L,R}$ and the real and imaginary parts of $\zeta_{L,R}$.
Note that the MZM case with $\zeta_{L,R}=0$ is characterized by $P$ and $Q$ remaining unchanged between the two parity sectors while $R$ changes sign. This corresponds to a $2\pi$ shift of $\phi$, which is the essential signature expected for a topological phase.